OpenStudy (anonymous):

Find the product of (x + 7)(x − 7). A) x^2 − 14x + 49 B) x^2 − 49 C) x^2 + 14x + 49 D) x^2 + 49

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

x^2-49

OpenStudy (dmezzullo):

@bahupepe explain ur answer plz

OpenStudy (anonymous):

exactly !

OpenStudy (nathan917):

The product for A. is 1) x^2 - 14x + 49 = (x-7) (x-7) = (x-7)^2

OpenStudy (nathan917):

For B. its X^2-49=0 Simple and best practice solution for X^2-49=0 equation. Check how easy it is, and learn it for the future. Our solution is simple, and easy to understand, so don`t hesitate to use it as a solution of your homework. If it's not what You are looking for type in the equation solver your own equation and let us solve it. Equation: Solution for X^2-49=0 equation: Simplifying X2 + -49 = 0 Reorder the terms: -49 + X2 = 0 Solving -49 + X2 = 0 Solving for variable 'X'. Move all terms containing X to the left, all other terms to the right. Add '49' to each side of the equation. -49 + 49 + X2 = 0 + 49 Combine like terms: -49 + 49 = 0 0 + X2 = 0 + 49 X2 = 0 + 49 Combine like terms: 0 + 49 = 49 X2 = 49 Simplifying X2 = 49 Take the square root of each side: X = {-7, 7}

OpenStudy (anonymous):

7x7=49 x.x=x^2 (+)(-)= - (x+7)(x-7) x.x-7x+7x-(7)(7) x^2-49